# Discretion in the Loop: Human Expertise in Algorithm-Assisted College Advising

### v2

@@ 6.3 Functional and Normative Roles of Human Discretion in Institutional Decision Making
Our study provides suggestive (though not confirmatory) evidence that students achieve better educational outcomes due to advisors’ discretionary contributions. This finding contributes to a long-standing debate across decision-making domains about whether human discretion helps or hinders the delivery of high-quality interventions, and when its institutional role should be increased, maintained or diminished in light of automated decision systems.
@@
Evidence across domains suggests that discretion is a double-edged sword. On the positive side, discretion enables flexible, “street-level” judgments by human decision-makers, allowing them to override algorithmic recommendations based on rich, real-time information to meet evolving goals [3, 31, 37, 62]. Discretion could also promote equity; for example, in child maltreatment call screening, call workers have helped close racial disparities by adjusting how they respond to algorithmic risk scores, mindful that the data driving these tools may reflect systemic biases [15, 24]. In addition to these benefits, our study suggests that advisor discretion acts as a bridge between predictions (e.g., risk scores) and effective interventions [44]. By eliciting and responding to non-algorithmic information from students, advisors select and deliver meaningful interventions that go beyond the capabilities of an algorithm. Whether advisor discretion also provides additional equity benefits—such as counterbalancing systemic biases of risk scores—remains an important question for future research.
@@
On the negative side, discretion can hinder decision making, particularly in settings where decision quality is thought to depend mostly on prediction accuracy [38] (though this is not our setting). In some predictive contexts, algorithms may outperform human forecasters [18] and so lead to improvements in decision quality. Even when human decision-makers have access to information beyond what the algorithm does [4], studies show that people often misjudge when to override or defer to algorithmic advice, leading to over-reliance or under-reliance [7]. Moreover, discretion can introduce bias, as demonstrated by the case of a pretrial risk assessment tool in Kentucky, where judges were more likely to override low-risk scores for Black defendants than for white defendants [2].
@@
Importantly, questions about discretion are not only about effectiveness but also about institutional values. The use of proprietary, opaque algorithms without a human in the loop can create accountability gaps, limiting the ability of students or advisors to contest decisions [13]. In other domains, it has been shown that rigid institutional mandates to follow algorithmic recommendations can undermine worker autonomy, generate resentment, and shift the balance of power away from human workers [62, 12]. Finally, those classified as “high-risk” by risk prediction tools may face stigmatization [69], and may express strong preferences for discretion over algorithmic decision making [35]. Particularly in educational contexts like advising, where personal connection and trust are critical, designing algorithmic systems must involve careful reflection on when and how to preserve human discretion.
@@ A.1.1 Baseline Characteristics
  • The student’s expected financial contribution for the 2016-2017 academic year (in twelve \$$\$1050$ intervals, ranging from \$$\$0-\$1050$ to \$$\$22,051$ or more)
@@ A.1.3 Meeting Logs
  • What initiated the advising interaction with the student (one option selected from: registration or program review, advising system alert, regularly scheduled advising, early alert or instructor initiated, and student intiated)
@@
The list of twenty interventions included:
@@
    Advised to seek academic support
@@
    Advised to correct schedule
@@
    Advised to declare or change major
@@
    Advised about progression issues
@@
    Advised to adjust course balance
@@
    Advised to see instructor
@@
    Advised to seek academic coaching
@@
    Advised to seek departmental or faculty guidance
@@
    Advised to address financial issues
@@
    Advised to address non-financial collegiate requirements
@@
    Advised to seek career counseling and service
@@
    Advised to seek engagement opportunities on campus
@@
    Advised to see professional counselor
@@
    Advised to seek disability services on campus
@@
    Advised to see medical professional
@@
    Advised to schedule an appointment
@@
    Advised about recommended schedule
@@
    Advised to see primary program advisor
@@
    Advised through the in-person development or review of the academic map/degree planner
@@ A.1.4 Data Quality
Figure 2: Normalized distributions of applied interventions per advisor, demonstrating an especially low diversity of interventions applied by Advisors 1 and 2. Advisors 4 and 5 are excluded due to the limited number of available comments for each.
There was some inconsistency in the way the advisors documented the interventions they performed during these meetings. In some cases, while there is clear evidence in the free-form comments that advisors performed a certain intervention during an interaction, they failed to mark the corresponding intervention boxes on the provided checklist. Advisor 1 in particular consistently selected limited sets of interventions (only ever indicating they had applied 10 of the 20 total interventions and only indicating they had applied four of them more than 100 times), even when their comments indicated that they had applied additional interventions. We show how the distribution of interventions applied varies for each advisor in Figure 2.
@@ A.2 Theoretical Framework
We prove Proposition 3.1 by defining two SCMs $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ such that both have the same joint distribution over $(A_{t},Y_{t+1})$ but only $\mathcal{M}_{1}$ meets the definition for expert targeting. Without loss of generality, we construct $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ such that there is no algorithmically available context $X_{t}$, corresponding to a setting where the algorithmic recommendations are not followed by the human decision-maker. [^28] We note the example can be modified to include $X_{t}$, for example by distributing it as a Bernoulli and adding a term to the equations for $Y_{t+1}$ and $A_{t}$ multiplying the entire expression by $X_{t}$ in both models.
@@
First, we show that $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are observationally equivalent. Since the SCM’s probability distributions are defined over the noise variables $\{\epsilon_{V}:V\in\mathbf{V}\}$, Table 4 demonstrates how the exogenous noise settings lead to different settings over model variables $\mathbf{V}$.
To derive the observational distribution $\mathbb{P}(A_{t},Y_{t+1})$ for each model, we sum the probabilities of all exogenous variable configurations that result in each setting of $A_{t}$ and $Y_{t+1}$. Table 5 presents the distribution for both models. Since the results are identical, this shows the models are observationally equivalent with respect to $\mathbb{P}(A_{t},Y_{t+1})$.
@@
Next, we show both of these models meet our assumption that SCMs are causally minimal. Causal minimality is equivalent to that, for all causal parents of $Y$, $X\in\mathbf{PA}_{Y}$, $X\not\perp Y\mid\mathbf{PA}_{Y}\backslash\{X\}$ [53, p.109]. Accordingly, for $\mathcal{M}_{1}$, we show that both $(1)Y_{t+1}\not\perp U_{t}|A_{t}$ and $(2)Y_{t+1}\not\perp A_{t}|U_{t}$. All probabilities are calculated by summing over the relevant rows of Table 4.
  1. 1.
    $Y_{t+1}\not\perp U_{t}|A_{t}$
    $\mathbb{P}(Y_{t+1}=1,U_{t}=0|A_{t}=1)$ $=\frac{\mathbb{P}(Y_{t+1}=1,U_{t}=0,A_{t}=1)}{\mathbb{P}(A_{t}=1)}$
    $=\frac{p^{2}-p^{3}}{2p-p^{2}}$
    $=\frac{p^{2}-p}{p-2}$
    $\mathbb{P}(Y_{t+1}=1|A_{t}=1)\times\mathbb{P}(U_{t}=0|A_{t}=1)$ $=\frac{\mathbb{P}(Y_{t+1}=1,A_{t}=1)}{\mathbb{P}(A_{t}=1)}\times\frac{\mathbb{P}(U_{t}=0,A_{t}=1)}{\mathbb{P}(A_{t}=1)}$
    $=\frac{-p^{3}+p^{2}+p}{2p-p^{2}}\times\frac{p-p^{2}}{2p-p^{2}}$
    $=\frac{p^{3}-2p^{2}+1}{p^{2}-4p+4}$
    Unless $p=1$ (ruled out by assumption), $\frac{p^{2}-p}{p-2}\neq\frac{p^{3}-2p^{2}+1}{p^{2}-4p+4}$. Thus, $\mathbb{P}(Y_{t+1}=1,U_{t}=0|A_{t}=1)\neq\mathbb{P}(Y_{t+1}=1|A_{t}=1)\times\mathbb{P}(U_{t}=0|A_{t}=1)$ and so $Y_{t+1}\not\perp U_{t}|A_{t}$.
  2. 2.
    $Y_{t+1}\not\perp A_{t}|U_{t}$
    $\mathbb{P}(Y_{t+1}=1,A_{t}=1|U_{t}=0)$ $=\frac{\mathbb{P}(Y_{t+1}=1,A_{t}=1,U_{t}=0)}{\mathbb{P}(U_{t}=0)}$
    $=\frac{p^{2}-p^{3}}{1-p}$
    $=p^{2}$
    $\mathbb{P}(Y_{t+1}=1|U_{t}=0)\times\mathbb{P}(A_{t}=1|U_{t}=0)$ $=\frac{p^{2}-p^{3}}{1-p}\times\frac{p-p^{2}}{1-p}$
    $=p^{2}\times p$
    $=p^{3}$
    Unless $p=0$ or $p=1$ (ruled out by assumption), $p^{2}\neq p^{3}$. Thus $\mathbb{P}(Y_{t+1}=1,A_{t}=1|U_{t}=0)\neq\mathbb{P}(Y_{t+1}=1|U_{t}=0)\times\mathbb{P}(A_{t}=1|U_{t}=0)$ and so $Y_{t+1}\not\perp A_{t}|U_{t}$.
Meanwhile, $\mathcal{M}_{2}$ meets causal minimality vacuously. Since there are only two model variables, each variable has (at most) one parent.
@@
    $=\frac{\mathbb{P}(Y_{t+1}=1,A_{t}=1,U_{t}=1)}{\mathbb{P}(U_{t}=1,A_{t}=1)}$ Definition of conditional prob.
    $=\frac{p}{p}$ Table 4 and simplification
    $=1$
@@
    $=\frac{\mathbb{P}(Y_{t+1}=1,A_{t}=0,U_{t}=1)}{\mathbb{P}(A_{t}=0,U_{t}=1)}$ Table 4 and simplification
    $=\mathbb{P}(Y_{t+1}|A_{t}=0,U_{t}=1)$ Definition of conditional prob.
@@
    Parents do/see is that conditioning and intervening on parent values of a variable are interchangeable [9, p.32]. It is true for all DAGs generated by SCMs [9, p.33].
@@
    $=\frac{\mathbb{P}(A_{t}=1,U_{t}=1}{\mathbb{P}(U_{t}=1)}$
    $=\frac{p}{p}$ Table 4 and simplification
    $=1$
@@
    $=\frac{p-p^{2}}{1-p}$ Algebraic manipulation
    $=\frac{\mathbb{P}(A_{t}=1,U_{t}=0)}{\mathbb{P}(U_{t}=0)}$ Table 4 and simplification
    $=\mathbb{P}(A_{t}=1|do(U_{t}=0))$ Definition of conditional prob. and parents do/see
@@
    Heterogeneous action:
    $\mathbb{P}(Y_{t}=1|do(A_{t}=1,U_{t}=1))$
    $-\mathbb{P}(Y_{t}=1|do(A_{t}=0,U_{t}=1))$ $=\mathbb{P}(Y_{t}=1|A_{t}=1,U_{t}=1)$
    $-\mathbb{P}(Y_{t}=1|A_{t}=0,U_{t}=1)$ Parents do/see
    $=\frac{\mathbb{P}(Y_{t}=1,A_{t}=1,U_{t}=1)}{\mathbb{P}(A_{t}=1,U_{t}=1)}$ Definition of conditional prob.
    $-\frac{\mathbb{P}(Y_{t}=1,A_{t}=0,U_{t}=1)}{\mathbb{P}(A_{t}=1,U_{t}=1)}$ Definition of conditional prob.
    $=\frac{p}{p}-\frac{0}{p}$
    $=1$
@@
    $=\frac{p^{2}-p^{3}}{p-p^{2}}-0$ Algebraic manipulation
    $=\frac{\mathbb{P}(Y_{t+1}=1,A_{t}=1,U_{t}=0)}{\mathbb{P}(A_{t}=1,U_{t}=0)}$
    $-\frac{\mathbb{P}(Y_{t+1}=1,A_{t}=0,U_{t}=0)}{\mathbb{P}(A_{t}=0,U_{t}=0)}$ Table 4 and simplification
    $=\mathbb{P}(Y_{t}=1|do(A_{t}=1,U_{t}=0))$
    $-\mathbb{P}(Y_{t}=1|do(A_{t}=0,U_{t}=0))$ Definition of conditional prob. and parents do/see
Meanwhile, note that $\mathcal{M}_{2}$ cannot meet targeted action nor heterogeneous action, since $U_{t}$ is not included in the model.
Thus, $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are causally minimal models inducing equivalent observational distributions $\mathbb{P}(A_{t},Y_{t+1})$, yet only $\mathcal{M}_{1}$ exhibits expertly targeted action. ∎
A.2.2 Proof of Proposition 3.3
Proof.
By the Markov condition, if $A_{t}\not\perp Y_{t+1}|\ \mathbf{X}_{t}$ then $A_{t}\not\perp_{\mathcal{G}_{\mathcal{M}}}Y_{t+1}|\ \mathbf{X}_{t}$. Thus, by the definition of d-separation, either $\mathbf{U}_{t}\to A_{t}$ or $A_{t}\to Y_{t+1}$.
First, consider the case that $A_{t}\rightarrow Y_{t+1}$. Then, for some $X_{t}=x$, $A_{t}=a$, $U_{t}=u$ and $Y_{t+1}=y$ we have:
$\mathbb{P}(Y_{t+1}=1,A_{t}=a|X_{t}=x,U_{t}=u)$ $\neq\mathbb{P}(Y_{t+1}=1|X_{t}=x,U_{t}=u)$
×$\quad\times\mathbb{P}(A_{t}=a|X_{t}=x,U_{t}=u)$ Causal minimality
$\mathbb{P}(Y_{t+1}=1|A_{t}=a,X_{t}=x,U_{t}=u)$
×$\quad\times\mathbb{P}(A_{t}=a|X_{t}=x,U_{t}=u)$ $\neq\mathbb{P}(Y_{t+1}=1|X_{t}=x,U_{t}=u)$
×$\quad\times\mathbb{P}(A_{t}=a|X_{t}=x,U_{t}=u)$ Product rule
$\mathbb{P}(Y_{t+1}=1|A_{t}=a,X_{t}=x,U_{t}=u)$ $\neq\mathbb{P}(Y_{t+1}=1|X_{t}=x,U_{t}=u)$ Cancellation
$\mathbb{P}(Y_{t+1}=y|A_{t}=a,X_{t}=x,U_{t}=u)$ $\neq\mathbb{P}(A_{t}=1)\times\mathbb{P}(Y_{t+1}=y|A_{t}=1,X_{t}=x,U_{t}=u)$
$\quad+\mathbb{P}(A_{t}=0)\times\mathbb{P}(Y_{t+1}=y|A_{t}=0,X_{t}=x,U_{t}=u)$ Law of total probability
$\mathbb{P}(Y_{t+1}=y|A_{t}=a,X_{t}=x,U_{t}=u)$ $\neq\mathbb{P}(Y_{t+1}=y|A_{t}=0,X_{t}=x,U_{t}=u)$∗Simplification
$\mathbb{P}(Y_{t+1}=y|do(A_{t}=1),X_{t}=x,U_{t}=u)$ $\neq\mathbb{P}(Y_{t+1}=y|do(A_{t}=0),X_{t}=x,U_{t}=u)$∗∗Parents do/see
$\mathbb{P}(Y_{t+1}=1|do(A_{t}=1),X_{t}=x,U_{t}=u)$ $\neq\mathbb{P}(Y_{t+1}=1|do(A_{t}=0),X_{t}=x,U_{t}=u)$ $Y_{t+1}$ binary
Recall that causal minimality is equivalent to that, for all causal parents of $Y$, $X\in\mathbf{PA}_{Y}$, $X\not\perp Y\mid\mathbf{PA}_{Y}\backslash\{X\}$ [53, p.109].
∗∗ Suppose $a=1$. Then by subtraction from both sides,
$\mathbb{P}(A_{t}=0)\times\mathbb{P}(Y_{t+1}=y|A_{t}=1,X_{t}=x,U_{t}=u)\neq\mathbb{P}(A_{t}=0)\times\mathbb{P}(Y_{t+1}=y|A_{t}=0,X_{t}=x,U_{t}=u)$
And so by cancellation,
$\mathbb{P}(Y_{t+1}=y|A_{t}=1,X_{t}=x,U_{t}=u)\neq\mathbb{P}(Y_{t+1}=y|A_{t}=0,X_{t}=x,U_{t}=u)$
The reasoning is symmetrical if $a=0$.
∗∗∗ Recall parents do/see is that conditioning and intervening on parent values of a variable are interchangeable [9, p.32]. It is true for all DAGs generated by SCMs [9, p.33].
Thus $A_{t}\rightarrow Y_{t+1}$ implies impactful action is met. $U_{t}\rightarrow A_{t}$ implies non-algorithmic action is met by a very similar argument. Thus, since either $A_{t}\rightarrow Y_{t+1}$ or $U_{t}\rightarrow A_{t}$ or both, also either impactful action is met, non-algorithmic action is met, or both.
[…]

### v3

@@ 7.4 The Practical and Normative Role of Discretion
Our finding that discretion plays a positive role in algorithm-assisted college advising contributes to two long-standing debates about algorithmic decision making: First, does human discretion help or hinder the delivery of high-quality interventions? And second, when should its institutional role be increased, maintained or diminished in light of automated decision systems?
@@
The empirical evidence on the first question is split [88]. On the positive side, discretion can enable flexible, “street-level” judgments by human decision makers, allowing them to override algorithmic recommendations based on real-time information and evolving goals [5, 40, 47, 79]. Discretion can also promote equity; for example, in child maltreatment call screening, call workers have helped close racial disparities by adjusting how they respond to algorithmic risk scores to counteract systemic biases in the data [20, 34]. In addition to these benefits, our study suggests that advisor discretion acts as a bridge between predictions (e.g., risk scores) and interventions [54]. By responding to non-algorithmic information from students, advisors deliver meaningful interventions that go beyond the capabilities of an algorithm. Whether advisor discretion also provides additional equity benefits—such as counterbalancing systemic biases of risk scores—remains an important question for future research.
@@
On the negative side, discretion can hinder decision making, particularly in settings where decision quality is thought to depend mostly on prediction accuracy [48] (though this is not our setting), as algorithms often outperform human forecasters [24]. Additionally, while human decision makers may have access to information beyond what the algorithm does [6], people may misjudge when to override or defer to algorithmic advice, leading to over-reliance or under-reliance [10]. Moreover, discretion can introduce bias, as demonstrated by the case of a pretrial risk assessment tool in Kentucky, where judges were more likely to override low-risk scores for Black defendants than for white ones [4].
@@
Ultimately, decisions about discretion cannot only be about effectiveness but also must incorporate institutional values. The use of proprietary, opaque algorithms without a human in the loop can create accountability gaps, limiting the ability of students or advisors to contest decisions [18]. It has been shown that rigid institutional mandates to follow algorithmic recommendations can undermine worker autonomy, generate resentment, and shift the balance of power away from human workers [79, 17]. Those classified as “high-risk” by risk prediction tools may face stigmatization [87], especially when predictions use demographic variables. Finally, students may prefer human discretion over algorithmic decision making [46], trusting trust human actors over algorithmic systems to prioritize their best interests [23]. Particularly in educational contexts like advising, where personal connection and trust are critical, designing algorithmic systems must involve careful reflection on when and how to preserve human discretion.
@@ A.1.1 Baseline Characteristics
  • The student’s expected financial contribution for the 2016-2017 academic year (in twelve $\mathdollar 1050$ intervals, ranging from $\mathdollar 0-\mathdollar 1050$ to $\mathdollar 22,051$ or more)
@@ A.1.3 Meeting Logs
  • What initiated the advising interaction with the student (one option selected from: registration or program review, advising system alert, regularly scheduled advising, early alert or instructor initiated, and student initiated)
@@
The list of twenty interventions were as follows, with descriptions from the MAAPS implementation guide provided to institutions participating in the study:
@@
    Advised to seek academic support: The student is encouraged to seek academic support hrough tutoring or supplemental instruction. The purpose of the intervention is to help the student earn a grade that will support their academic progression and prevent them from failing or dropping a course.
@@
    Advised to correct schedule: The student is encouraged to seek academic support through tutoring or supplemental instruction. The purpose of the intervention is to help the student earn a grade that will support their academic progression and prevent them from failing or dropping a course.
@@
    Advised to declare or change major: Student is encouraged to declare or change their major based on advisor recommendations (potentially based on predictive analytics). Purpose of the intervention is to ensure student is able to complete the academic markers needed to fulfill all degree requirements in a timely manner.
@@
    Advised about progression issues: The purpose of the intervention is to encourage the student to address curricular progression issues. Examples of such issues include enrollment in low number of hours, dropping classes, withdrawing from classes, receiving incompletes for coursework and other issues.
@@
    Advised to adjust course balance: The purpose of the intervention is to ensure that the student takes the appropriate course balance and does not take a disproportionate number of high risk courses within the same term. The appropriate course balance information may be based on predictive analytics or the grades that the student has earned in prerequisite or gateway courses.
@@
    Advised to see instructor: The student is encouraged to seek the advice/assistance of their course instructor in response to early alert notification for poor grades, lack of attendance or other academic concerns.
@@
    Advised to seek academic coaching: The student is encouraged to seek specialized assistance related to non-cognitive academic skills including time management, note taking, reading text books, motivation etc.
@@
    Advised to seek departmental or faculty guidance: The purpose of the intervention is to encourage the student to build relationships with faculty in order to better understand professional, career or graduate school opportunities related to the student’s academic program. Use this category when the student is advised to see the faculty for reasons other than academic performance in the classroom. Please use category 6 to record advised to see faculty based on academic concerns.
@@
    Advised to address financial issues: The purpose of the intervention is to encourage the student to address financial issues that undermine retention or academic progression. This would include issues related to FAFSA and FAFSA completion, Student Account balances, scholarships etc.
@@
    Advised to address non-financial collegiate requirements: Purpose of intervention is to encourage student to address non-financial collegiate requirements that undermine retention or academic progression. Requirements may include testing and placement, document submission and maybe related to: vaccinations and health, final transcripts, housing information etc.
@@
    Advised to seek career counseling and service: The purpose of intervention is to encourage student to seek advice about how to develop career and professional plan.
@@
    Advised to seek engagement opportunities on campus: The purpose of intervention is to encourage the student to get involved in co-curricular, recreation or wellness oriented activities.
@@
    Advised to see professional counselor: The purpose of intervention is to encourage the student to see a professional counselor for non-academic personal issues.
@@
    Advised to seek disability services on campus: The purpose of intervention is to encourage the student to seek support of campus disability services because of learning disability or physical handicap.
@@
    Advised to see medical professional: The purpose of intervention is to encourage the student to see a medical professional in response to health or wellness-related concerns.
@@
    Advised to schedule an appointment: The purpose of intervention is to encourage the student to schedule a follow-up appointment with their MAAPS advisor.
@@
    Advised about recommended schedule: The purpose of the intervention is to recommend courses that student should take in the future. This intervention should be recorded when the MAAPS advisor provides a recommended schedule. If the MAAPS advisors correct a schedule, the intervention should be recorded as 2, 4 or 5 above.
@@
    Advised to see primary program advisor: The purpose of the intervention is to encourage the student to meet with a program advisor familiar with academic and non-academic program requirements and opportunities including internships, seminars, and other activities related to the student’s program of study.
@@
    Advised through the in-person development or review of the academic map/degree planner: This category should be indicated if an advisor works 1-on-1 with the student to develop or review an academic map or planner.
@@ A.1.4 Data Quality
Figure 3: Histograms of applied interventions per advisor, demonstrating an especially low diversity of interventions applied by Advisor 1. Advisors 4 and 5 are excluded due to the limited number of available comments for each.
Our exploratory qualitative analysis reveals some inconsistency in the way that one advisor, Advisor 1, documented the interventions they performed during meetings. While sometimes there is clear evidence in the free-form comments that Advisor 1 performed a certain intervention during an interaction, they failed to mark the corresponding intervention boxes on the provided checklist. Additionally, Advisor 1 only indicated they had applied 10 of the 20 total interventions and only indicated they had applied four of them more than 100 times, even when their comments indicated that they had applied many additional interventions. We show how the distribution of interventions applied varies for each advisor in Figure 3. To account for reduced intervention data quality, we repeat the $\text{ExpertTest}^{*}$ excluding Advisor 1 to establish robustness of the results.
@@ A.2 Theoretical Framework
We prove Proposition 4.1 by defining two SCMs $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ such that both have the same joint distribution over $(A_{t},Y_{t+1})$ but only $\mathcal{M}_{1}$ meets the definition for expertly targeted action. Without loss of generality, we construct $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ such that there is no algorithmically available context $X_{t}$, corresponding to a setting where the algorithmic recommendations are ignored by the human decision-maker. [^28] We note the example can be modified to include $X_{t}$, for example by distributing it as a Bernoulli and adding a term to the equations for $Y_{t+1}$ and $A_{t}$ multiplying the entire expression by $X_{t}$ in both models.
@@
First, we show that $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are observationally equivalent. To derive the observational distribution $\mathbb{P}(A_{t},Y_{t+1})$ for each model, we sum the probabilities of all exogenous variable configurations that result in each setting of $A_{t}$ and $Y_{t+1}$. Table 5 demonstrates how the exogenous noise settings lead to different settings over model variables $\mathbf{V}$. Table 6 presents the observational distribution over $A_{t}$ and $Y_{t+1}$ for both models. Since the results are identical, this shows the models are observationally equivalent with respect to $\mathbb{P}(A_{t},Y_{t+1})$.
@@
Next, we show both of these models meet Assumptions 4.1 through 4.4. By construction, both models meet Assumption 4.1, inducing DAGs meeting the graphical constraints, and Assumption 4.2, as noise variables are distributed independently. We now show that both $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ meet Assumption 4.4 of faithfulness, from which causal minimality (Assumption 4.3) follows. Demonstrating faithfulness requires showing that, for all variables, a lack of $d$-separation implies statistical dependence. We show this is true for $\mathcal{M}_{1}$ below, with probabilities calculated using Table 5.
  1. (1)
    $U_{t}\not\perp A_{t}$
    $\mathbb{P}(U_{t}=0,A_{t}=1)$ $=p-p^{2}$
    $\mathbb{P}(U_{t}=0)\times\mathbb{P}(A_{t}=1)$ $=p^{3}-3p^{2}+2p$
    For all $\in(0,1), pp$$p^{2}\neq p^{3}-3p^{2}+2p$. Thus, $U_{t}\not\perp A_{t}$.
  2. (2)
    $U_{t}\not\perp Y_{t+1}$
    $\mathbb{P}(U_{t}=0,Y_{t+1}=1)$ $=p^{2}-p^{3}$
    $\mathbb{P}(U_{t}=0)\times\mathbb{P}(Y_{t+1}=1)$ $=p^{4}-2p^{3}+p$
    For all $p\in(0,1)$, $p^{2}-p^{3}\neq p^{4}-2p^{3}+p$. Thus, $U_{t}\not\perp Y_{t+1}$.
  3. (3)
    $A_{t}\not\perp Y_{t+1}$
    $\mathbb{P}(A_{t}=0,Y_{t+1}=1)$ $=0$
    $\mathbb{P}(A_{t}=0)\times\mathbb{P}(Y_{t+1}=1)$ $=-p^{5}+3p^{4}-2p^{3}-p^{2}$
    For all $p\in(0,1)$, $p^{5}+3p^{4}-2p^{3}-p^{2}\neq 0$. Thus, $A_{t}\not\perp Y_{t+1}$.
  4. (4)
    $U_{t}\not\perp A_{t}|Y_{t+1}$
    $\mathbb{P}(U_{t}=1,A_{t}=1|Y_{t+1}=1)$ $=\frac{p^{2}}{-p^{2}+p+1}$
    $\mathbb{P}(U_{t}=1|Y_{t+1}=1)\times\mathbb{P}(A_{t}=1|Y_{t+1}=1)$ $=\frac{1}{-p^{2}+p+1}$
    For all $p\in(0,1)$, $\frac{p^{2}}{-p^{2}+p+1}\neq\frac{1}{-p^{2}+p+1}$. Thus, $U_{t}\not\perp A_{t}|Y_{t+1}$.
  5. (5)
    $U_{t+1}\not\perp Y_{t+1}|A_{t}$
    $\mathbb{P}(Y_{t+1}=1,U_{t}=0|A_{t}=1)$ $=\frac{p^{2}-p^{3}}{2p-p^{2}}$
    $\mathbb{P}(Y_{t+1}=1|A_{t}=1)\times\mathbb{P}(U_{t}=0|A_{t}=1)$ $=\frac{p^{3}-2p^{2}+1}{p^{2}-4p+4}$
    For all $p\in(0,1)$, $\frac{p^{2}-p}{p-2}\neq\frac{p^{3}-2p^{2}+1}{p^{2}-4p+4}$. Thus, $Y_{t+1}\not\perp U_{t}|A_{t}$.
  6. (6)
    $A_{t}\not\perp Y_{t+1}|U_{t}$
    $\mathbb{P}(Y_{t+1}=1,A_{t}=1|U_{t}=0)$ $=p^{2}$
    $\mathbb{P}(Y_{t+1}=1|U_{t}=0)\times\mathbb{P}(A_{t}=1|U_{t}=0)$ $=p^{3}$
    For all $p\in(0,1)$, $p^{2}\neq p^{3}$. Thus, $Y_{t+1}\not\perp A_{t}|U_{t}$.
To satisfy faithfulness, $\mathcal{M}_{2}$ must have $A_{t}\not\perp Y_{t+1}$. Since $\mathcal{M}_{1}$ is observationally equivalent to $\mathcal{M}_{2}$ over $(A_{t},Y_{t+1})$, this follows from (3).
@@
    $=1$ Calculated from Table 5
@@
    $=\mathbb{P}(Y_{t+1}|A_{t}=0,U_{t}=1)$ Calculated from Table 5
@@
    Parents do/see is that conditioning and intervening on parent values of a variable are interchangeable [13, p.32]. It is true for all DAGs generated by SCMs [13, p.33].
@@
    $=1$ Calculated from Table 5
@@
    $=\mathbb{P}(A_{t}=1|U_{t}=0)$ Calculated from Table 5
    $=\mathbb{P}(A_{t}=1|do(U_{t}=0))$ Parents do/see
@@
    Heterogeneous effects:
    $\Delta(1)$ $=\mathbb{P}(Y_{t}=1|A_{t}=1,U_{t}=1)-\mathbb{P}(Y_{t}=1|A_{t}=0,U_{t}=1)$ Parents do/see
    $=1$ Calculated from Table 5
@@
    $=\mathbb{P}(Y_{t}=1|A_{t}=1,U_{t}=0)-\mathbb{P}(Y_{t}=1|A_{t}=0,U_{t}=0))$ Calculated from Table 5
    $=\Delta(0)$ Parents do/see
Meanwhile, $\mathcal{M}_{2}$ can meet neither of targeted action nor heterogeneous effects, since $U_{t}$ is not included in the model. Thus, $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ induce identical observational distributions $\mathbb{P}(A_{t},Y_{t+1})$, yet only $\mathcal{M}_{1}$ exhibits expertly targeted action. ∎
$\mathbf{U}_{t}$$\mathbf{X}_{t}$$A_{t}$$Y_{t+1}$